OpenStudy (anonymous):

Given the functions f(x) = x2 + 6x − 1, g(x) = –x2 + 2, and h(x) = 2x2 − 4x + 3, rank them from least to greatest based on their axis of symmetry. f(x), g(x), h(x) h(x), g(x), f(x) g(x), h(x), f(x) h(x), f(x), g(x)

OpenStudy (campbell_st):

do you know how to find the axis of symmetry..?

OpenStudy (anonymous):

not when it is in standard form

OpenStudy (karatechopper):

You would need to convert to vertex form right? @campbell_st

OpenStudy (anonymous):

nvm -b --- 2a right?

OpenStudy (campbell_st):

ok... so if an parabola is in the form \[y = ax^2 + bx + c\] the axis of symmtery is \[x = \frac{-b}{2a}\] this may be familar as its part of the general quadratic formula so on the 1st question you have a = 1 and b = 6 so the axis of symmetry is \[x = \frac{-6}{2 \times 1}\] just solve it hope that makes sense

OpenStudy (campbell_st):

then repeat the process for the 2nd and 3rd equations.

OpenStudy (anonymous):

so fgh? A

OpenStudy (campbell_st):

makes sense to me

OpenStudy (anonymous):

not A